A jet completes a journet of 1600km in 1 1/2 hours.
what is the average speed of the jet
how long would it have taken to travel the first 300km

Answer by checkley79(3341)

D=RT
1,600=R*1.5
R=1,600/1.5
R=1,066.67 KMH. ANS.
300=1,066.67T
T=300/1,066.67
T=.28 HOURS. OR 16.8 MINUTES.